# Tail Distribution of Regret in Optimistic Reinforcement Learning

### v1

@@ 5  Proof Analysis
$\leq$ $\exp\left(-\frac{(x-\lceil K^{\gamma}\rceil H-m_{K})^{2}}{2H^{3}K}\right).$
Proof of Theorem 2.
@@
$\mathbbm{E}[R_{K}]=$ $\mathbbm{E}[R_{K}^{(0)}]+\mathbbm{E}[R_{K}^{(1)}+R_{K}^{(2)}]$
$\leq$ $H\lceil K^{\gamma}\rceil+\mathbbm{E}\left[(R_{K}^{(1)}+R_{K}^{(2)})\left(\mathbbm{1}\left\{\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right\}+\mathbbm{1}\left\{\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right\}\right)\right]$
$=$ $H\lceil K^{\gamma}\rceil+\mathbbm{E}\left[(R_{K}^{(1)}+R_{K}^{(2)})\mathbbm{1}\left\{\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right\}\right]+\mathbbm{E}\left[(R_{K}^{(1)}+R_{K}^{(2)})\mathbbm{1}\left\{\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right\}\right]$
$\leq$ $H\lceil K^{\gamma}\rceil+\mathbbm{E}\left[(R_{K}^{(1)}+R_{K}^{(2)})\mathbbm{1}\left\{\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right\}\right]+H(K-\lceil K^{\gamma}\rceil)\mathbb{P}\left(\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right).$ (6)
Under the event $\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}$, we have that $R_{K}^{(1)}<0$ and that
$R_{K}^{(2)}\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}2b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\eta_{k,h+1}^{(2)}(s_{h+1}^{k}).$
Using the bounds from the proof of Theorem 1, we obtain
$\mathbbm{E}\left[(R_{K}^{(1)}+R_{K}^{(2)})\mathbbm{1}\left\{\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right\}\right]$
$\quad\leq m_{K}+\mathbbm{E}\left[\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\eta_{k,h+1}^{(2)}(s_{h+1}^{k})\right)\mathbbm{1}\left\{\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right\}\right]$
$\quad=m_{K}+\mathbbm{E}\left[\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\eta_{k,h+1}^{(2)}(s_{h+1}^{k})\right)\right]+\mathbbm{E}\left[\left(\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\eta_{k,h+1}^{(2)}(s_{h+1}^{k})\right)\mathbbm{1}\left\{\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right\}\right]$
$\quad=m_{K}+\mathbbm{E}\left[\left(\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\eta_{k,h+1}^{(2)}(s_{h+1}^{k})\right)\mathbbm{1}\left\{\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right\}\right]$
$\quad\leq m_{K}+H^{2}(K-\lceil K^{\gamma}\rceil)\mathbb{P}\left(\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right)$ (7)
where $m_{K}$ is as in (5) and the last equality follows by the fact that $\{\eta_{k,h+1}^{(2)}(s_{h+1}^{k})\}$ is a Martingale difference sequence.
@@
$\mathbb{E}[R_{K}]\leq H\lceil K^{\gamma}\rceil+m_{K}+2H^{2}(K-\lceil K^{\gamma}\rceil)\mathbb{P}\left(\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right)$
By the analysis of the proof of Theorem 1, the probability of the event $\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}$ is bounded as follows:
$\mathbb{P}\left(\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right)\leq\delta_{K}=SAH\begin{cases}\displaystyle\ K\exp\left(-2\mu K^{\alpha}\right),&\text{{KD}},\\[8.61108pt] \displaystyle K\exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}$

### v2

@@ 5.1 Proof of Theorem 1
$=\delta_{K}^{\mathrm{mb}}(\gamma)+\mathbb{P}\left(R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\geq x-\lceil K^{\gamma}\rceil H\right).$ (since $x>\lceil K^{\gamma}\rceil H$)
We proceed with bounding the accumulative regret $R_{K}^{(1)}=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}_{k}}(s_{0}^{k})\right)$, under the event $\bar{\mathcal{G}}^{K}$.
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (5))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\mathbbm{1}\{\mathcal{G}^{k}\}$
$\qquad\qquad\qquad\qquad-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\mathbbm{1}\{\mathcal{G}^{k}\}$
$\qquad\qquad\qquad\qquad-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ ($V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})=Q_{h}^{\bm{\pi}^{k}}(s_{h}^{k},a_{h}^{k})\leq Q_{h}^{*}(s_{h}^{k},a_{h}^{k})$)
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (6))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\},$ (7)
where $\tilde{\Delta}_{h}^{k}(s_{h}^{k})\coloneqq V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})$. We express $\tilde{\Delta}_{h}^{k}(s_{h}^{k})$ recursively, introducing terms that can be bounded either via concentration inequalities for $\hat{\bm{P}}_{k}$ around $\bm{P}^{*}$ or by the bonus functions.
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $\leq\mathbbm{1}\{\mathcal{G}^{k}\}b_{h}^{k}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}\left(\hat{P}^{k}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})-P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\right)V_{h+1}^{k}(s^{\prime})$
$\qquad+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\left(V_{h+1}^{k}(s^{\prime})-V_{h+1}^{\bm{\pi}^{k}}(s^{\prime})\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mathbbm{1}\{\mathcal{G}^{k}\}b_{h}^{k}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\frac{H-h-1}{2}\mathbbm{1}\{\mathcal{G}^{k}\}\|P_{h}^{*}(\cdot\mid s_{h}^{k},a_{h}^{k})-\hat{P}_{h}^{k}(\cdot\mid s_{h}^{k},a_{h}^{k})\|_{1}$
$\qquad+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\tilde{\Delta}_{h+1}^{k}(s^{\prime})$
$\leq 2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{s^{\prime}}P^{*}_{h}(s^{\prime}\mid s_{h}^{k},a_{h}^{k})\tilde{\Delta}_{h+1}^{k}(s^{\prime})$
$=2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\mathbb{E}\left[\mathbbm{1}\{\mathcal{G}^{k}\}\tilde{\Delta}_{h+1}^{k}(s_{h+1}^{k})\mid\mathcal{F}_{h}^{k}\right]$
$=2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\tilde{\Delta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}-\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}.$ (8)
where $(a)$ follows from Hölder’s inequality, and noting that $\sum_{s}(P_{h}^{*}(s\mid s_{h}^{k},a_{h}^{k})-\hat{P}_{h}^{k}(s\mid s_{h}^{k},a_{h}^{k}))=0$, and the sequence $\{\tilde{\eta}^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\tilde{\eta}^{k}_{h}(s_{h}^{k})=\tilde{\Delta}^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\tilde{\Delta}^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right],$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $\tilde{\eta}^{k}_{H}\equiv 0$ for all $k\in[K]$. Summing (8) over $\{h,h+1,\cdots,H-1\}$, the terms telescope, yielding
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\leq 2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))-\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k}).$ (9)
Next, the claim the following inequality holds:
$\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}$
$\qquad\leq\,\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k}).$ (10)
If $a_{h}^{k}\in\mathcal{A}_{h}^{*}(s_{h}^{k})$, the inequality holds trivially. Suppose that $a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$. Note that,
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $=(V^{k}_{h}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}$
$\geq(V^{*}_{h}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}$ (by $\mathbbm{1}\{\mathcal{G}^{k}\}$ and $V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\leq Q_{h}^{*}(s_{h}^{k},a_{h}^{k})$)
$=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$
$\geq\mathrm{gap}^{*}\mathbbm{1}\{\mathcal{G}^{k}\}$ ($a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$)
which implies $\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}=\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)$ given $a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})$. Now, we have the following cases:
  1. 1.
    If $\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})\geq 2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))$, then
    $\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)$ $=\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$
    $\leq-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})$ (by (9))
  2. 2.
    If $2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))>-\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})$, then
    $\mathrm{gap}^{*}\mathbbm{1}\{\mathcal{G}^{k}\}\leq\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}$ $\leq 4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k})),$ (by (9))
    which implies
    $\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)=4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k})).$
Substituting (10) into (7) yields
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ $\leq\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(4\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{l=h}^{H-1}b^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$
$\qquad\qquad-2\mathbbm{1}\{\mathcal{G}^{k}\}\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$\leq\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathrm{clip}\left(4Hb^{k}_{l}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\right)$ (by (4))
$\qquad\qquad-2\sum_{h=0}^{H-1}\sum_{l=h}^{H-1}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\tilde{\eta}_{l+1}^{k}(s_{l+1}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\,\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$
$\qquad\qquad-\sum_{h=0}^{H-1}\left(2\left(\sum_{l=0}^{h}\mathbbm{1}\{a_{l}^{k}\notin\mathcal{A}_{l}^{*}(s_{l}^{k})\}\right)\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})+\eta^{k}_{h+1}(s_{h+1}^{k})\right)\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\,\left[\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (11)
where the sequence $\{\bar{\eta}_{h}^{k}\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\bar{\eta}^{k}_{h+1}=2\left(\sum_{l=0}^{h}\mathbbm{1}\{a_{l}^{k}\notin\mathcal{A}_{l}^{*}(s_{l}^{k})\}\right)\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})+\eta^{k}_{h+1}(s_{h+1}^{k}),$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$, and satisfies $|\bar{\eta}^{k}_{h+1}|\leq 5H^{2}$.
Next, we bound the first term in (11). We have
$\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$ $\leq H\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}\mathrm{clip}\left(4Hb^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\right)$
$\leq H\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}\mathrm{clip}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n_{h}^{k}(s_{h}^{k},a_{h}^{k})\}}}\right)$
$\leq H\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}\sum_{n=1}^{\infty}\mathrm{clip}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n\}}}\right)$
$=H\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}\sum_{n=1}^{\bar{n}}\left(4H^{2}\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n\}}}\right)$
$\leq 8H^{4}SA\sqrt{0.5S\ln 2+\mu K^{\alpha}}\sqrt{\bar{n}}$
$\leq\frac{32H^{6}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}}\eqqcolon m_{K}^{\mathrm{mb}}$ (12)
where
$\bar{n}=\left\lfloor\frac{16H^{4}(0.5S\ln 2+\mu K^{\alpha})}{{\mathrm{gap}^{*}}^{2}}\right\rfloor.$
Combining the bounds, we have
$\mathbb{P}\left(\left\{R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\geq x-\lceil K^{\gamma}\rceil H\right\}\right)$
$\qquad\qquad\leq\mathbb{P}\left(\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\left(m_{K}^{\mathrm{mb}}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)\geq x-\lceil K^{\gamma}\rceil H\right)$
$\qquad\qquad\leq\mathbb{P}\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\geq x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}}\right)$ (since $x>\lceil K^{\gamma}\rceil H$)
$\qquad\qquad\leq\exp\left(-\frac{(x-\lceil K^{\gamma}\rceil H-m_{K}^{\mathrm{mb}})^{2}}{50H^{5}K}\right).$
where the last inequality follows by applying Azuma–Hoeffding.
Proof of Proposition 1.
@@
$\mathbb{E}[R_{K}]=$ $\mathbb{E}[R_{K}^{(0)}]+\mathbb{E}[R_{K}^{(1)}]$
$\leq$ $H\lceil K^{\gamma}\rceil+\mathbb{E}\left[R_{K}^{(1)}\left(\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}+\mathbbm{1}\left\{\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right\}\right)\right]$
$=$ $H\lceil K^{\gamma}\rceil+\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]+\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\left(\bar{\mathcal{G}}^{K}\right)^{c}\right\}\right]$
$\leq$ $H\lceil K^{\gamma}\rceil+\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]+H(K-\lceil K^{\gamma}\rceil)\mathbb{P}\left(\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right).$ (13)
Note that by (11) and (12), we have
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\leq\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\left(m_{K}^{\mathrm{mb}}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)$
Hence, we obtain
$\mathbb{E}\left[R_{K}^{(1)}\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]$ $\leq m_{K}^{\mathrm{mb}}+\mathbb{E}\left[\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)\mathbbm{1}\left\{\bar{\mathcal{G}}^{K}\right\}\right]$
$=m_{K}^{\mathrm{mb}}+\mathbb{E}\left[-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right]+\mathbb{E}\left[\left(\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{\eta}_{h+1}^{k}\right)\mathbbm{1}\left\{\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right\}\right]$
$\leq m_{K}^{\mathrm{mb}}+5KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma).$ (14)
@@
$\mathbb{E}[R_{K}]\leq H\lceil K^{\gamma}\rceil+m_{K}^{\mathrm{mb}}+6KH^{3}\delta_{K}^{\mathrm{mb}}(\gamma).$